Martina leaves home at 9 A.M., bicycling at a rate of
24 mi/h. Two hours later, John leaves, driving at the rate of 48 mi/h. At what time
will John catch up with Martina?

Answer by ankor@dixie-net.com(12692)

Martina leaves home at 9 A.M., bicycling at a rate of 24 mi/h. Two hours
later, John leaves, driving at the rate of 48 mi/h. At what time will John
catch up with Martina?
:
One thing to remember about these "catch-up" problems is when it occurs, both
parties will have traveled the same distance. Make a distance equation>
:
Let t = M's time in hrs from 9 AM
Then (t-2) = J's time when he overtakes M
:
Dist = speed*time
:
J's dist = M's dist
48(t-2) = 24t
48t - 96 = 24t
48t - 24t = + 96
24t = 96
t = 96/24
t = 4 hrs from 9 AM would be 1 pm when J catches up
:
:
Check our solution using the dist equation:
48(2) = 24(4)
:
Did this make sense to you??